North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size

Q6 - 17 November 2021

Ian reports a fine mild day at Grosmont on Wdnesday with Bill Dobson, Steve Hyman, Ed Bolam, Nigel Hall, Nigel Bill, Jon Bradley, Gordon Wells, David Potter and Ian Pearson.

That morning, the Q6 was moved from No 2 road onto the pit outside Deviation Shed. This allowed Ian Pearson to work underneath to adjust the oil feed pipes into a position where the oil drops centrally along the top of the horn guides. Bill Dobson and David Potter pressure tested the atomisers and found them to be OK with no leakage from joints and castings. These are intended to be fitted in the New Year. Afterwards they concentrated on the new axle box lubricator, changing the thimbles, eight of which were made at Hopetown last week. Bit of a fiddly job, and only four were fitted this week.

Nigel Bill arrived with a new battery for the defibrillator which was fitted and is now functioning again. Nigel then set about sorting out and tidying up our cabinet drawers upstairs, resulting in some rubbish being thrown out. Steve Hyman is in the process of making a plate around the lefthand tender water valve. This has been cut to size and a stability bracket was made but it requires welding under the new plate next week. Gordon Wells came in and repaired the outstanding PAT testing failures. These were sent to Hopetown for re-testing on Thursday 18th.

Jon Bradley arrived and finished the new access gate, but it does need a step fitting to the front.

Later, Nigel Hall and Ed Bolam lagged the steam heat pipe after an earlier complaint from a fireman who had nearly received a burn while disposing of the ashes from the ash pan. The pipe area lagged was from the front damper forwards about 6 feet.

In the afternoon, David Potter had his site safety induction with Paul Whickham.